OpenStudy (anonymous):

Subtract the polynomials: (6x3 - 4x + 5) - (2x3 - x2 + 7x - 10). Enter your answer as a polynomial in descending order below, and use the caret (^) for exponents.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Subtract the polynomials: (6x3 - 4x + 5) - (2x3 - x2 + 7x - 10). Enter your answer as a polynomial in descending order below, and use the caret (^) for exponents.

OpenStudy (anonymous):

combine like terms: 6x^3-2x^3= 4x^3 -4x-7x= -11x 5+10= 15 -(-x^2)= x^2 answer: 4x^3+x^2-11x+15

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

Find the sum: (x5 - 3x4 + x2 - 4) + (5x4 + 3x2 + 2x + 1). Enter your answer as a polynomial in descending order below, and use the caret (^) for exponents. HELP

OpenStudy (anonymous):

Find the sum: (x5 - 3x4 + x2 - 4) + (5x4 + 3x2 + 2x + 1). Enter your answer as a polynomial in descending order below, and use the caret (^) for exponents.